# Synergizing RAG and Reasoning: A Systematic Review

### v1

@@ 1. Introduction
Models like DeepSeek-R1, with strong reasoning capabilities, inspire new directions for RAG systems. As shown in Figure 1, recent research explores integrating formal reasoning frameworks with knowledge retrieval. This approach optimizes retrieval through logic-driven query reformulation and uses reasoning to analyze and validate retrieved knowledge, creating cognitive synergy between retrieval and generation. This paradigm aims to overcome conventional limitations, enabling intelligent systems with rigorous logic and reliable knowledge use. From the trend perspective, more and more methods are combining reasoning and retrieval abilities through reinforcement learning (RL), marking a new trend in the LRM era. Meanwhile, prompt-based approaches continue to rapidly evolve, with researchers aiming to achieve results through workflow design while keeping model parameters frozen. It is also noticeable that reliance solely on tuning methods is steadily decreasing, indicating that additional supervised fine-tuning (SFT) offers limited improvement at the current stage of development.
@@
1) From Ambiguous Semantic Matching to Logic-Driven Targeted Retrieval. Traditional RAG relies on semantic similarity for document retrieval, which is sensitive to phrasing variations. Advanced reasoning allows deep logical analysis of queries (e.g., causal links, conditional constraints) to dynamically refine retrieval strategies (Guan et al., 2025). For example, to answer ”How to reduce postoperative infection risks in diabetes patients?”, the system prioritizes retrieving ”blood glucose control thresholds” and ”antibiotic usage guidelines” over simply matching ”diabetes postoperative care.” This approach supports multi-hop retrieval by breaking down complex queries into sequential sub-queries while preserving cross-document coherence through reasoning chains.
@@
  • Systematic Taxonomy. A multi-dimensional framework is presented to systematically examine the objectives, paradigms, and methodologies for combining RAG with reasoning capabilities, establishing clear classification criteria across technical dimensions.
@@
  • Open Resource Platform [^1] https://openrag.notion.site/open-rag-base?pvs=4 Through the OpenRAG platform, we provide a rich and multi-dimensional review of related work. It is convenient for readers to quickly search and compare different methods.
@@ 2.1. Definition
The definition of reasoning in modern AI systems remains an evolving construct, particularly within the context of LRMs exemplified by DeepSeek R1 and OpenAI O1. Here, under the scope of LLM, we formalize reasoning as a structured, multi-step process that dynamically decomposes complex problems, generates intermediate hypotheses, and iteratively refines solutions through logical and evidence-based transformations. Mathematically, let a reasoning process $\mathcal{R}$ be defined as a tuple $\langle\mathcal{K}_{p},\mathcal{K}_{r},\mathcal{S}_{t},\Phi\rangle$, where $\mathcal{K}_{p}$ denotes parametric knowledge embeddings, $\mathcal{K}_{r}$ represents retrieved contextual knowledge, $\mathcal{S}_{t}=\{s_{0},s_{1},...,s_{n}\}$ constitutes the evolving state sequence with $s_{0}$ as the initial query and $s_{n}$ as the final response, and $\Phi:\mathcal{S}_{i}\times\mathcal{K}_{p}\times\mathcal{K}_{r}\rightarrow\mathcal{S}_{i+1}$ defines the state transition function.
@@
The reasoning process exhibits three defining characteristics. First, it is inherently multi-step, systematically decomposing complex problems into intermediate cognitive states (e.g., sub-question generation or temporary conclusions) rather than pursuing direct input-output mapping. Second, it generates novel (new) knowledge or fact—synthesizing implicit relationships, deriving latent constraints, or reformulating problems in ways not explicitly present in the initial input or parametric memory (e.g., transforming ”Is A greater than B?” into comparative subquestions about A and B’s attributes). Crucially, these representations are not merely retrieved but dynamically constructed through the reasoning trajectory. Third, the process is teleological—its architecture and termination conditions are explicitly optimized for complex problem resolution, where complexity is measured by the necessity of state transitions or the insufficiency of direct retrieval from either parametric ($\mathcal{K}_{p}$) or external ($\mathcal{K}_{r}$) knowledge sources. This stands in stark contrast to atomic inference, which lacks such deliberate state construction and goal-aware iteration.
@@
The distinction between reasoning and inference manifests most saliently in their computational signatures. While inference $\mathcal{I}$ constitutes a single-step conditional probability computation $P(y|x)=\prod_{t=1}^{T}P(y_{t}|x,y_{<t})$, reasoning $\mathcal{R}$ implements a meta-process coordinating multiple inference calls through explicit state management $\mathcal{R}(x)=\Phi_{1}\circ\Phi_{2}\circ\cdots\circ\Phi_{n}(x)$ This multi-phase architecture enables systematic error correction through backtracking mechanisms and dynamic retrieval refinement, features fundamentally absent in conventional inference pipelines. The operational boundary emerges when state transitions involve explicit symbolic manipulation (equation restructuring in mathematical reasoning) or knowledge graph traversal (temporal reasoning over retrieved events), distinguishing true reasoning from mere multi-step inference.
@@
(Better Reasoning) §3.2, text width=3.5cm [ ActiveRAG (Xu et al., 2024); AgenticReasoning (Wu et al., 2025b); CoRAG (Wang et al., 2025a); CR-planner (Li et al., 2024a); DeepRAG (Guan et al., 2025); Deepsolution (Li et al., 2025b); KBQA-O1 (Luo et al., 2025); OpenRAG (Islam et al., 2024); PIKE (Wang et al., 2025b); R1-Seacher (Song et al., 2025); RAG-Gym (Xiong et al., 2025b); ReARTeR (Sun et al., 2025); ReSearch (Chen et al., 2025b); MedRAG (Zhao et al., 2025); RARE (Tran et al., 2024); RetroRAG (Xiao et al., 2025); KG-RAR (Wu et al., 2025a); RetrievalPRM (Zhu et al., 2025b); MRD-RAG (Chen et al., 2025d); Search-O1 (Li et al., 2025a); StePO-Rec (Bi et al., 2025a), text width=10.0cm, node options=align=left ] ] ] [Paradigms §4, text width=2.5cm, for tree=fill=orange!20,font=[Pre-defined Workflow § 4.1, text width=2.8cm [Reasoning in Pre-Retrieval,text width=2.1cm [ LeReT (Hsu et al., 2024); O1-Embedder (Yan et al., 2025); planRAG (Lee et al., 2024); UAR (Cheng et al., 2024); MMOA-RAG (Chen et al., 2025e); MedRAG (Zhao et al., 2025); FRAG (Gao et al., 2025); Insight-RAG (Pezeshkpour and Hruschka, 2025); ChainRAG (Zhu et al., 2025a); RetrievalPRM (Zhu et al., 2025b); REAPER (Joshi et al., 2024); AdaptiveRAG (Jeong et al., 2024), text width=8.2cm, node options=align=left ] ] [Reasoning in Post-Retrieval,text width=2.1cm [ ActiveRAG (Xu et al., 2024); ARM (Chen et al., 2025f); MRD-RAG (Chen et al., 2025d); FG-RAG (Hong et al., 2025); ToG2 (Ma et al., 2024), text width=8.2cm, node options=align=left ] ] [Hybrid Reasoning,text width=2.1cm [ IR-COT (Trivedi et al., 2022a); FinSearch (Li et al., 2024c); ITER-RETGEN (Shao et al., 2023); LevelRAG (Zhang et al., 2025); HiRAG (Zhang et al., 2024); MetaRAG (Zhou et al., 2024); RetroRAG (Xiao et al., 2025); KG-RAR (Wu et al., 2025a); DeepNote (Wang et al., 2024e), text width=8.2cm, node options=align=left ] ] ] [Dynamic Workflow §4.2, text width=2.8cm [Proactivity-Driven Reasoning,text width=2.2cm [ AgenticReasoning (Wu et al., 2025b); DeepRAG (Guan et al., 2025); CoRAG (Wang et al., 2025a); Co-STORM (Jiang et al., 2024a); PIKE (Wang et al., 2025b); Search-O1 (Li et al., 2025a); R1-Searcher (Song et al., 2025), text width=8.1cm, node options=align=left ] ] [Reflection-Driven Reasoning,text width=2.2cm [ Flare (Jiang et al., 2023); OpenRAG (Islam et al., 2024); WriteHere (Xiong et al., 2025a); ReaRAG (Lee et al., 2025); Self-RAG (Asai et al., 2023), text width=8.1cm, node options=align=left ] ] [Feedback-Driven Reasoning,text width=2.2cm [ SmartRAG (Gao et al., 2024a); CR-Planner (Li et al., 2024a); MCTS-KBQA (Xiong et al., 2025c); DeepSolution (Li et al., 2025b); RAG-Gym (Xiong et al., 2025b); ReZero (Dao and Le, 2025); ReSearch (Chen et al., 2025b); RARE (Tran et al., 2024); DeepResearcher (Zheng et al., 2025); PORAG (Srinivas and Runkana, 2025); MCTS-RAG (Hu et al., 2025); KBQA-O1 (Luo et al., 2025), text width=8.1cm, node options=align=left ] ] ] ] [Implementation §5, text width=2.5cm, for tree=fill=blue!20,font=[Resoning Method §5.1, text width=2.8cm, [LLM/CoT, text width=2.0cm [ ActiveRAG (Xu et al., 2024); AdaptiveRAG (Jeong et al., 2024); O1-Embedder (Yan et al., 2025); DeepNote (Wang et al., 2024e); HiRAG (Zhang et al., 2024); MetaRAG (Zhou et al., 2024); MMOA-RAG (Chen et al., 2025e); RetroRAG (Xiao et al., 2025); PORAG (Srinivas and Runkana, 2025); KG-RAR (Wu et al., 2025a); MRD-RAG (Chen et al., 2025d); PlanRAG (Lee et al., 2024), text width=8.2cm, node options=align=left ] ] [Special Token Prediction, text width=2.0cm [ Self-RAG (Asai et al., 2023); SmartRAG (Gao et al., 2024a); OpenRAG (Islam et al., 2024); R1-Searcher (Song et al., 2025); ReZero (Dao and Le, 2025); ReSearch (Chen et al., 2025b); ReaRAG (Lee et al., 2025); Search-O1 (Li et al., 2025a), text width=8.2cm, node options=align=left ] ] [Search-Driven Reasoning, text width=2.0cm [ OminiThink (Xi et al., 2025); DeepRAG (Guan et al., 2025); CoRAG (Wang et al., 2025a); DeepSolution (Li et al., 2025b); ReARTeR (Sun et al., 2025); KBQA-O1 (Luo et al., 2025); WriteHere (Xiong et al., 2025a); RARE (Tran et al., 2024); MCTS-KBQA (Xiong et al., 2025c); StePO-Rec (Bi et al., 2025a), text width=8.2cm, node options=align=left ] ] [Reasoning on Graph, text width=2.0cm [ FinSearch (Li et al., 2024c); ToG2 (Ma et al., 2024); LighRAG (Guo et al., 2024); FRAG (Gao et al., 2025); FG-RAG (Hong et al., 2025); MedRAG (Zhao et al., 2025), text width=8.2cm, node options=align=left ] ] [External Solver,text width=2.0cm [ ARM (Chen et al., 2025f), text width=8.2cm, node options=align=left ] ] ] [Optimization §5.2, text width=2.8cm, for tree=fill=blue!20 [Prompt-Based, text width=2.0cm [ Co-STORM (Jiang et al., 2024a); Agentic Reasoning (Wu et al., 2025b); FinSearch (Li et al., 2024c); PlanRAG (Lee et al., 2024); HiRAG (Zhang et al., 2024); MetaRAG (Zhou et al., 2024); MedRAG (Zhao et al., 2025); RARE (Tran et al., 2024); ReaRAG (Lee et al., 2025); RetroRAG (Xiao et al., 2025); FRAG (Gao et al., 2025); KG-RAR (Wu et al., 2025a); MRD-RAG (Chen et al., 2025d); FG-RAG (Hong et al., 2025); MCTS-RAG (Hu et al., 2025); DeepSolution (Li et al., 2025b); StePO-Rec (Bi et al., 2025a), text width=8.2cm, node options=align=left ] ] [Tuning-Based, text width=2.0cm [ KBQA-Q1 (Luo et al., 2025); O1-Embedder (Yan et al., 2025); DeepRAG (Guan et al., 2025); CoRAG (Wang et al., 2025a); MCTS-KBQA (Xiong et al., 2025c); RetrievalPRM (Zhu et al., 2025b); REAPER (Joshi et al., 2024); UAR (Cheng et al., 2024), text width=8.2cm, node options=align=left ] ] [RL-Based, text width=2.0cm [PIKE (Wang et al., 2025b); LeReT (Hsu et al., 2024); RAG-Gym (Xiong et al., 2025b); ReARTeR (Sun et al., 2025); SmartRAG (Gao et al., 2024a); CR-Planner (Li et al., 2024a); DeepRetrieval (Jiang, 2025); DeepNote (Wang et al., 2024e); MMOA-RAG (Chen et al., 2025e); ReZero (Dao and Le, 2025); ReSearch (Chen et al., 2025b); DeepResearcher (Zheng et al., 2025); PORAG (Srinivas and Runkana, 2025); R1-Search (Song et al., 2025), text width=8.2cm, node options=align=left ] ] ] ] ]

### v2

@@ 1. Introduction
Models like DeepSeek-R1, with strong reasoning capabilities, inspire new directions for RAG systems. As shown in Figure 1, recent research explores integrating formal reasoning frameworks with knowledge retrieval. This approach optimizes retrieval through logic-driven query reformulation and uses reasoning to analyze and validate retrieved knowledge, creating cognitive synergy between retrieval and generation. This paradigm aims to overcome conventional limitations, enabling intelligent systems with rigorous logic and reliable knowledge use. From a trend perspective, an increasing number of methods combine reasoning and retrieval abilities through reinforcement learning (RL), marking a new direction in the LRM era. Meanwhile, prompt-based approaches continue to rapidly evolve, with researchers aiming to achieve results through workflow design while keeping model parameters frozen. Notably, sole reliance on tuning methods is steadily decreasing, suggesting limited improvements from additional fine-tuning at this stage.
@@
1) From Ambiguous Semantic Matching to Logic-Driven Targeted Retrieval. Traditional RAG relies on semantic similarity for retrieval; however, it is sensitive to phrasing variations. Advanced reasoning allows deep logical analysis of queries (e.g., causal links, conditional constraints) to dynamically refine retrieval strategies (Guan et al., 2025). For example, to answer ”How to reduce postoperative infection risks in diabetes patients?”, the system prioritizes retrieving ”blood glucose control thresholds” and ”antibiotic usage guidelines” over simply matching ”diabetes postoperative care”. This approach supports multi-hop retrieval by breaking down complex queries into sequential sub-queries while preserving cross-document coherence through reasoning chains.
@@
  • Systematic Taxonomy. We present a multi-dimensional framework to systematically examine the objectives, paradigms, and methodologies for combining RAG with reasoning capabilities, establishing clear classification criteria across technical dimensions.
@@
  • Open Resource Platform [^1] https://openrag.notion.site/open-rag-base?pvs=4 Through the OpenRAG platform, we provide a rich, multi-dimensional review of related work, which allows readers to quickly search and compare different methods.
@@ 2.1. Definition
The definition of reasoning in modern AI systems remains an evolving construct, particularly within the context of LRMs exemplified by DeepSeek R1 and OpenAI O1. Here, under the scope of LLMs, we formalize reasoning as a structured, multi-step process that dynamically decomposes complex problems, generates intermediate hypotheses, and iteratively refines solutions through logical and evidence-based transformations. Mathematically, let a reasoning process $\mathcal{R}$ be defined as a tuple $\langle\mathcal{K}_{p},\mathcal{K}_{r},\mathcal{S}_{t},\Phi\rangle$, where $\mathcal{K}_{p}$ denotes parametric knowledge embeddings, $\mathcal{K}_{r}$ represents retrieved contextual knowledge, $\mathcal{S}_{t}=\{s_{0},s_{1},\ldots,s_{n}\}$ constitutes the evolving state sequence with $s_{0}$ as the initial query and $s_{n}$ as the final response, and $\Phi:\mathcal{S}_{i}\times\mathcal{K}_{p}\times\mathcal{K}_{r}\rightarrow\mathcal{S}_{i+1}$ defines the state transition function.
@@
The reasoning process exhibits three defining characteristics. First, it is inherently multi-step, systematically decomposing complex problems into intermediate cognitive states (e.g., sub-question generation or temporary conclusions) rather than pursuing direct input-output mapping. Second, it generates novel knowledge or facts — synthesizing implicit relationships, deriving latent constraints, or reformulating problems in ways not explicitly present in the initial input or parametric memory (e.g., transforming ”Is A greater than B?” into comparative subquestions about A and B’s attributes). Crucially, these representations are not merely retrieved but dynamically constructed through the reasoning trajectory. Third, the process is teleological — its architecture and termination conditions are explicitly optimized for complex problem resolution, where complexity is measured by the necessity of state transitions or the insufficiency of direct retrieval from either parametric ($\mathcal{K}_{p}$) or external ($\mathcal{K}_{r}$) knowledge sources. This stands in stark contrast to atomic inference, which lacks such deliberate state construction and goal-aware iteration.
@@
The distinction between reasoning and inference manifests most saliently in their computational signatures. While inference $\mathcal{I}$ constitutes a single-step conditional probability computation $P(y|x)=\prod_{t=1}^{T}P(y_{t}|x,y_{<t})$, reasoning $\mathcal{R}$ implements a meta-process coordinating multiple inference calls through explicit state management $\mathcal{R}(x)=\Phi_{1}\circ\Phi_{2}\circ\cdots\circ\Phi_{n}(x)$. This multi-phase architecture enables systematic error correction through backtracking mechanisms and dynamic retrieval refinement — features fundamentally absent in conventional inference pipelines. The operational boundary emerges when state transitions involve explicit symbolic manipulation (equation restructuring in mathematical reasoning) or knowledge graph traversal (temporal reasoning over retrieved events), distinguishing true reasoning from mere multi-step inference.
@@
(Better Reasoning) §3.2, text width=3.5cm [ ActiveRAG (Xu et al., 2024); AgenticReasoning (Wu et al., 2025b); CoRAG (Wang et al., 2025a); CR-planner (Li et al., 2024a); DeepRAG (Guan et al., 2025); Deepsolution (Li et al., 2025b); KBQA-O1 (Luo et al., 2025); OpenRAG (Islam et al., 2024); PIKE (Wang et al., 2025b); R1-Seacher (Song et al., 2025); RAG-Gym (Xiong et al., 2025b); ReARTeR (Sun et al., 2025); ReSearch (Chen et al., 2025b); MedRAG (Zhao et al., 2025); RARE (Tran et al., 2024); RARE(Wang et al., 2025d); RetroRAG (Xiao et al., 2025); KG-RAR (Wu et al., 2025a); RetrievalPRM (Zhu et al., 2025b); MRD-RAG (Chen et al., 2025d); Search-O1 (Li et al., 2025a); StePO-Rec (Bi et al., 2025a), text width=10.0cm, node options=align=left ] ] ] [Paradigms §4, text width=2.5cm, for tree=fill=orange!20,font=[Pre-defined Workflow § 4.1, text width=2.8cm [Reasoning in Pre-Retrieval,text width=2.1cm [ LeReT (Hsu et al., 2024); O1-Embedder (Yan et al., 2025); planRAG (Lee et al., 2024); UAR (Cheng et al., 2024); MMOA-RAG (Chen et al., 2025e); MedRAG (Zhao et al., 2025); FRAG (Gao et al., 2025); Insight-RAG (Pezeshkpour and Hruschka, 2025); ChainRAG (Zhu et al., 2025a); RetrievalPRM (Zhu et al., 2025b); REAPER (Joshi et al., 2024); AdaptiveRAG (Jeong et al., 2024), text width=8.2cm, node options=align=left ] ] [Reasoning in Post-Retrieval,text width=2.1cm [ ActiveRAG (Xu et al., 2024); ARM (Chen et al., 2025f); MRD-RAG (Chen et al., 2025d); FG-RAG (Hong et al., 2025); ToG2 (Ma et al., 2024), text width=8.2cm, node options=align=left ] ] [Hybrid Reasoning,text width=2.1cm [ IR-COT (Trivedi et al., 2022a); FinSearch (Li et al., 2024c); ITER-RETGEN (Shao et al., 2023); LevelRAG (Zhang et al., 2025); HiRAG (Zhang et al., 2024); MetaRAG (Zhou et al., 2024); RetroRAG (Xiao et al., 2025); KG-RAR (Wu et al., 2025a); DeepNote (Wang et al., 2024e), text width=8.2cm, node options=align=left ] ] ] [Dynamic Workflow §4.2, text width=2.8cm [Proactivity-Driven Reasoning,text width=2.2cm [ AgenticReasoning (Wu et al., 2025b); DeepRAG (Guan et al., 2025); CoRAG (Wang et al., 2025a); Co-STORM (Jiang et al., 2024a); PIKE (Wang et al., 2025b); Search-O1 (Li et al., 2025a); R1-Searcher (Song et al., 2025), text width=8.1cm, node options=align=left ] ] [Reflection-Driven Reasoning,text width=2.2cm [ Flare (Jiang et al., 2023); OpenRAG (Islam et al., 2024); WriteHere (Xiong et al., 2025a); ReaRAG (Lee et al., 2025); Self-RAG (Asai et al., 2023), text width=8.1cm, node options=align=left ] ] [Feedback-Driven Reasoning,text width=2.2cm [ SmartRAG (Gao et al., 2024a); CR-Planner (Li et al., 2024a); MCTS-KBQA (Xiong et al., 2025c); DeepSolution (Li et al., 2025b); RAG-Gym (Xiong et al., 2025b); ReZero (Dao and Le, 2025); ReSearch (Chen et al., 2025b); RARE (Tran et al., 2024); DeepResearcher (Zheng et al., 2025); PORAG (Srinivas and Runkana, 2025); MCTS-RAG (Hu et al., 2025); KBQA-O1 (Luo et al., 2025), text width=8.1cm, node options=align=left ] ] ] ] [Implementation §5, text width=2.5cm, for tree=fill=blue!20,font=[Resoning Method §5.1, text width=2.8cm, [LLM/CoT, text width=2.0cm [ ActiveRAG (Xu et al., 2024); AdaptiveRAG (Jeong et al., 2024); O1-Embedder (Yan et al., 2025); DeepNote (Wang et al., 2024e); HiRAG (Zhang et al., 2024); MetaRAG (Zhou et al., 2024); MMOA-RAG (Chen et al., 2025e); RetroRAG (Xiao et al., 2025); PORAG (Srinivas and Runkana, 2025); KG-RAR (Wu et al., 2025a); MRD-RAG (Chen et al., 2025d); PlanRAG (Lee et al., 2024), text width=8.2cm, node options=align=left ] ] [Special Token Prediction, text width=2.0cm [ Self-RAG (Asai et al., 2023); SmartRAG (Gao et al., 2024a); OpenRAG (Islam et al., 2024); R1-Searcher (Song et al., 2025); ReZero (Dao and Le, 2025); ReSearch (Chen et al., 2025b); ReaRAG (Lee et al., 2025); Search-O1 (Li et al., 2025a), text width=8.2cm, node options=align=left ] ] [Search-Driven Reasoning, text width=2.0cm [ OminiThink (Xi et al., 2025); DeepRAG (Guan et al., 2025); CoRAG (Wang et al., 2025a); DeepSolution (Li et al., 2025b); ReARTeR (Sun et al., 2025); KBQA-O1 (Luo et al., 2025); WriteHere (Xiong et al., 2025a); RARE (Tran et al., 2024); MCTS-KBQA (Xiong et al., 2025c); StePO-Rec (Bi et al., 2025a), text width=8.2cm, node options=align=left ] ] [Reasoning on Graph, text width=2.0cm [ FinSearch (Li et al., 2024c); ToG2 (Ma et al., 2024); LighRAG (Guo et al., 2024); FRAG (Gao et al., 2025); FG-RAG (Hong et al., 2025); MedRAG (Zhao et al., 2025), text width=8.2cm, node options=align=left ] ] [External Solver,text width=2.0cm [ ARM (Chen et al., 2025f), text width=8.2cm, node options=align=left ] ] ] [Optimization §5.2, text width=2.8cm, for tree=fill=blue!20 [Prompt-Based, text width=2.0cm [ Co-STORM (Jiang et al., 2024a); Agentic Reasoning (Wu et al., 2025b); FinSearch (Li et al., 2024c); PlanRAG (Lee et al., 2024); HiRAG (Zhang et al., 2024); MetaRAG (Zhou et al., 2024); MedRAG (Zhao et al., 2025); RARE (Tran et al., 2024); ReaRAG (Lee et al., 2025); RetroRAG (Xiao et al., 2025); FRAG (Gao et al., 2025); KG-RAR (Wu et al., 2025a); MRD-RAG (Chen et al., 2025d); FG-RAG (Hong et al., 2025); MCTS-RAG (Hu et al., 2025); DeepSolution (Li et al., 2025b); StePO-Rec (Bi et al., 2025a), text width=8.2cm, node options=align=left ] ] [Tuning-Based, text width=2.0cm [ KBQA-Q1 (Luo et al., 2025); O1-Embedder (Yan et al., 2025); DeepRAG (Guan et al., 2025); CoRAG (Wang et al., 2025a); MCTS-KBQA (Xiong et al., 2025c); RetrievalPRM (Zhu et al., 2025b); REAPER (Joshi et al., 2024); RARE(Wang et al., 2025d); UAR (Cheng et al., 2024), text width=8.2cm, node options=align=left ] ] [RL-Based, text width=2.0cm [PIKE (Wang et al., 2025b); LeReT (Hsu et al., 2024); RAG-Gym (Xiong et al., 2025b); ReARTeR (Sun et al., 2025); SmartRAG (Gao et al., 2024a); CR-Planner (Li et al., 2024a); DeepRetrieval (Jiang, 2025); DeepNote (Wang et al., 2024e); MMOA-RAG (Chen et al., 2025e); ReZero (Dao and Le, 2025); ReSearch (Chen et al., 2025b); DeepResearcher (Zheng et al., 2025); PORAG (Srinivas and Runkana, 2025); R1-Search (Song et al., 2025), text width=8.2cm, node options=align=left ] ] ] ] ]